• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 08:35:34 on November 22, 2009

    Entry number 301772

    keyword=Compton Cavity locks now

    Cavity is locking now with around 750 W of power.

    M1 H -3155 V -1285

    M2 H -1350 V -90

    the mirror might have moved when the cavity enclosure was opened.